How much must you deposit today to withdraw M9000 in 5years if simple interest of 5% is earned
Answer by addingup(3677)

A= Amount
C= Initial capital
r= interest rate
n= number of years
............
simple interest:
A= C(1+r)^n
................
9000= C(1+0.05)^5
9000= C(1.05)^5
9000= C(1.27) Divide both sides by 1.276
9000/1.276= C
C= 7053.29 is the initial amount you need